Question 6

  1. Table 10-5
    Ronald Cummings purchased 800 shares of Barnum Corp.'s common stock for $15 per share. Barnum Corp.'s shares have a par value of $2.


    Referring to Table 10-5, if Barnum Corp. declared and issued a two-for-one stock split, and later declared and paid on the same day a cash dividend of $1.00 per share, what journal entry would Ronald Cummings make in order to record the cash dividend from Barnum Corp.?

    Investment in the A1 Resort 1,600
    Gain from Stock Split 1,600

    Investment in the A1 Resort 3,200
    Cash 800
    Gain from Stock Split 3,200
    Dividend Income 800

    Cash 1,600
    Dividend Income 1,600

    Cash 1,600
    Dividend Income 800
    Gain from Stock Split 800

    Cash 800
    Dividend Income 800
Answer #1
Shares held by Ronald Cummings after two-for-one stock split = 800*2 = 1600
Dividends received = 1600*1 = $1600
Cash 1,600
      Dividend Income 1,600
Option C is correct